Bedoo manufactures circular garden ponds that are 1.3 feet deep and have a diameter of 8 feet.
They are made out of synthetic rubber liners which she buys in rolls at a cost of $4.20 per square foot.
What does the rubber cost Bedoo to manufacture each pond
1
Expert's answer
2011-12-09T09:53:03-0500
The pond is a cylinder without a bottom, so the surface area of a pond is
S = pi*R^2 + 2*pi*R*H, where R = D/2 = 4 feet.
So,
S = pi*R^2 + 2*pi*R*H = pi*4^2 + 2*pi*4*1.3 ≈ 82.9380 squared feet.
and the price of the rubber is
S/P = 82.9380/4.20 = $19.7471 per pond.